OpenStudy (anonymous):

Help find the x and y intercepts 3x + 5y = 0 5x + 3y = 16

OpenStudy (amistre64):

x y intercepts? or x and y that they have in common?

OpenStudy (anonymous):

What each have in common

OpenStudy (amistre64):

does it matter how you solve it?

OpenStudy (anonymous):

It is a system

OpenStudy (anonymous):

solved by addition

OpenStudy (amistre64):

3x + 5y = 0 [-5] 5x + 3y = 16 [3] -15x -25y = 0 15x + 9y = 48 ---------------- -16y = 48 y = -3

OpenStudy (amistre64):

i think addition is the same thing as elimination....right?

OpenStudy (anonymous):

What I did was this (-3)3x + 5y = 0 (5)5x + 3y = 16 -9x - 15x = 0 25x + 15x = 80 16/16 \[\times\] 80/16 I got x = 5

OpenStudy (amistre64):

3x + 5y = 0 [-3] 5x + 3y = 16 [5] -9x -15y = 0 25x +15y = 80 -------------- 16x = 80 x = 5

OpenStudy (anonymous):

Can I find Y by using 3x + 5y = 0 ?

OpenStudy (amistre64):

once you find a solid value for a variable; you can just use it in an equation to find the other...

OpenStudy (anonymous):

I got +3 when I used 3x + 5y = 0

OpenStudy (amistre64):

no you dint lol

OpenStudy (amistre64):

3x + 5y = 0 3(5) + 5y = 0 15 +5y = 0 5y = -15 y = -15/5 y = -3

OpenStudy (anonymous):

Whoops! I forgot to add the - sign! Thanks :D

OpenStudy (anonymous):

I needed to send you my question so you can understand it better Complete the sales receipt below by finding the prices, the subtotal, the sales tax, and the total. McCOY'S FURNITURE 1 @ 490.00 SOFA 2 @ 60.00 END TABLE 1 @ 360.00